FIVE TIMES A NUMBER IS 4 MORE THAN THE NUMBER

1.FIVE TIMES A NUMBER IS 4 MORE THAN THE NUMBER. FIND THE NUMBER.
2.THE DIFFERENCE BETWEEN TWO NUMBERS IS 10 AND THEIR SUM IS 24. FIND THE NUMBERS
Answer #1
1.FIVE TIMES A NUMBER IS 4 MORE THAN THE NUMBER. FIND THE NUMBER. 5x = x+4 4x=4 x=1 -------------- 2.THE DIFFERENCE BETWEEN TWO NUMBERS IS 10 AND THEIR SUM IS 24. FIND THE NUMBERS x-y=10 x+y=24 --------- Add the equations to get: 2x=34 x=17 Substitute to get: 17+y=24 y = 7 =========== Cheers, Stan H.
answered by: mk-tintin